OpenStudy (anonymous):

use fundamental identities to simplify the expression (sin theta/cos theta) + (cos theta/sin theta)

OpenStudy (mathstudent55):

\( \tan \theta = \dfrac{\sin \theta}{\cos \theta }\) \( \cot \theta = \dfrac{\cos \theta}{\sin \theta }\)

OpenStudy (anonymous):

tan theta+cot theta